A station emits 96MHz frequency at the rate of 100KW. Calculate the number of photons emitted per second.
1
Expert's answer
2019-06-10T12:36:55-0400

the number of photons is

"N = \frac{P}{\epsilon} = \frac{N}{h\nu} = \frac{10^{5}}{6.63\cdot96\cdot10^{-28}} = 1.57\cdot10^{30}"

units/s.